# Mass Shootings, Community Mobility, and the Relocation of Economic Activity

### v1

@@ Abstract
Using foot traffic data for over 150,000 points of interest (POIs) near the sites of 42 mass shootings (2018-2022, U.S.), we evaluate the spatial-temporal impact of the tragic events on community mobility and relocation of economic activities. Visits to nearby POIs decrease, while farther away POIs experience increased foot traffic, implying that communities shift their activities away from the shooting sites. The impact is stronger when stronger trauma responses are expected. Our results suggest that mass shootings drive significant displacements of economic activities and can consequently lead to welfare losses due to distortions in optimal choices of time and location.
@@ 1 Introduction
Mass shootings have a profound effect on survivors, witnesses, and communities that experience them. The substantial loss of innocent lives is tragic on its own and clearly demonstrates that these events come at a great cost. However, the broader literature in economics, psychology, and beyond suggests that loss of life does not constitute the sole effect of mass shootings. Rather, it is accompanied by a variety of consequences on various fronts, including: worsening mental health for both direct victims and the surrounding community (Lowe & Galea, 2017; Sharkey & Shen, 2021; Wozniak et al., 2020); exacerbation of political polarization (Donohue et al., 2019; Luca et al., 2020; Yousaf, 2021); and decreases in school enrollment and performance (Beland & Kim, 2016; Levine & McKnight, 2024; Bharadwaj et al., 2021).
@@
Although these tragic events are the subject of a growing number of papers, evaluations of their economic consequences are rare. To the best of our knowledge, Brodeur & Yousaf (2022) is the only published study that provides a comprehensive analysis of mass shootings from an economic perspective. Using annual data, the authors find that mass shootings reduce the number of jobs and establishments in targeted counties by about 1.8% and 1.3%, respectively. They also find detrimental effects on earnings and housing prices in the years following the shootings.
@@
In contrast to Brodeur & Yousaf (2022), which presents year-over-year and county-level effects, our analysis is centered on a more immediate and direct economic outcome: changes in visits to locations within five miles of a shooting site, right after the incident. More specifically, we answer the following questions: Do mass shootings result in the displacement of foot traffic toward locations farther from the shooting sites? Do mobility changes differ by location characteristics, such as substitutability? And are these changes related to the perceived risk of victimization or traumatic responses to the event?
@@
For our analysis, we use the mass shootings that occurred in the United States between 2018 and 2022, as documented by Mother Jones. This list excludes shootings from more conventionally motivated crimes, such as robbery or gang activities, and thus allows our analysis to be grounded on non-targeted and unpredictable events. To measure mobility, we use Safegraph’s “weekly patterns” dataset, which includes foot traffic data and relevant characteristics for points of interest (POIs). Then, we employ event study and Difference-in-Differences (DiD) regressions to estimate the causal impact of mass shootings on community mobility over a period spanning ten weeks before and twenty weeks after the incidents.
@@
Our empirical findings can be summarized as follows. First, visits to nearby POIs decrease significantly after a mass shooting, while farther away POIs experience increased foot traffic, suggesting a relocation of economic activity. This pattern tends to amplify overtime, leading to a significant gap in weekly visits between regions. Second, POIs that can be substituted with relative ease due to their business nature or the presence of a large number of competitors in the region exhibit more intensive changes. POIs that are difficult to find alternatives, such as daycare facilities, show no significant effect. [^1] We discuss the welfare implications of the relocation of economic activities in Section 4.1. Even without measurable changes, welfare can still be affected. If the preference for the current location decreases due to a mass shooting but relocation is infeasible because of limited alternatives, a welfare loss may occur. Finally, the mobility changes and relocation of economic activity are stronger when the shooting occurs in a previously safer area, when it receives more media coverage, and for POIs that are more heavily visited before the incident. Combined, these results suggest that people attempt to conduct their activities away from the locations marked by these tragic events, especially under the conditions where stronger trauma responses are expected.
@@
This paper contributes to the literature in several ways. First, it sheds light on an important driver for the long-term, macroeconomic effects of mass shootings found in other studies (e.g., Brodeur & Yousaf, 2022). After an event of this type, communities relocate economic activities away from the tragic site, likely incurring friction costs or efficiency losses due to distortions in optimal location choices. This can ultimately worsen various macro-level outcomes, including wages, employment rates, and housing prices. Second, this study contributes to the broader literature on mass shootings. Our results reveal behavioral changes in response to trauma, linking the economic consequences to those documented in studies on trauma psychology (Wozniak et al., 2020; Sharkey & Shen, 2021), the economics of crime (Braakman, 2012; Dustmann & Fasani, 2014), criminology and sociology (Garofalo, 1981; Hale, 1996), and public health (Marquet et al., 2020; Guite et al., 2006). Finally, this study finds that recreational spaces, educational facilities, grocery stores, and places for temporary accommodation are among the most impacted locations. This opens up an opportunity for future studies to use the variations in activities in these categories to identify other important relationships between economic variables.
@@ 2.2 Community Mobility
The foot traffic data used in this study comes from Safegraph’s weekly patterns dataset, which is based on a collection of cellphone location data from a large list of applications that send pings - that is, requests for GPS data - to the phone in which they are installed. [^3] Papers using the Safegraph data include Chen & Rohla (2018), Chen et al. (2022), Chen & Pope (2020). Safegraph aggregates pings to produce a large panel-type dataset containing the number of weekly visits to distinct points of interest (POIs). [^4] See https://www.safegraph.com/guides/visit-attribution-white-paper for more details on Safegraph’s methodology In the dataset, a point of interest is any physical place of a non-residential nature, such as a restaurant, a museum, or a church, and it can belong to any of 180+ industry groups. In addition to tracking the number of visits to individual POIs within specified date ranges, the data include information that allows us to identify the locations and types of POIs: unique ID, latitude and longitude, location name, and North American Industry Classification System (NAICS) code.
Our main analysis uses a total of 4.2 million observations for about 153 thousand distinct POIs located within five miles of the mass shooting sites, covering a period of 30 weeks for each event - 10 weeks before the shooting and 20 weeks after. The five-mile radius was chosen based on nonparametric estimation results suggesting that the impact of mass shootings on mobility vanishes beyond five miles from the shooting site. [^5] Online Appendix B.2 contains the nonparametric results.
Six shooting events took place in more than one location. These events, known as sprees, require a special handling since our analysis relies heavily on the distance between a POI and its associated shooting site. In these cases, an overall centroid may be used to represent the crime scenes, but we have opted to include all crime scenes as separate shooting sites, including all POIs within a five mile radius of each of the crime scenes, and assigning the minimum distance-to-shooting-site in the event of overlaps. This strategy improves precision and is also more adequate in our analysis given that, in some cases, the distance between crime scenes is substantial (e.g., the Atlanta Massage Parlor Shootings’ victims were 30 miles apart). In online Appendix A, events are individually identified as being sprees or having truncated time series. [^6] The Safegraph data start in January 2018 and end in December 2022, so the mobility data are truncated for some shooting events. For instance, the Marjorie Stoneman Douglas High School shooting, which occurred on February 14th, 2018, does not have a full set of 10 weeks of data before the event, and it is thus truncated on the left. Seven out of the 42 mass shootings have such truncated time series.
@@ 2.4 Descriptive Analysis
(b) Evolution of Aggregate Visits and Active POIs
(c) Evolution of Aggregate Visits and Active POIs - By Distance Band
@@
*Note: Subfigures (b) and (c) plot the evolution of visits and active POIs, using the week immediately before the shooting as the baseline. Sprees and events with incomplete time series are excluded.
@@
This section offers an introductory description of the mobility patterns observed in the data. Figure 1(a) shows the changes in weekly visits to McDonald’s locations in Chicago after the Mercy Hospital shooting on November 19th, 2018.888This exercise uses a franchise to minimize heterogeneity across locations. The y-axis represents the changes in weekly visits at each location compared to the week right before the shooting. Clearly, most McDonald’s locations experienced a decrease in weekly visits in both week 1 and week 8 after the shooting (red dots), but we do see some locations farther from the shooting site experienced an increase in visits in the latter period (green dots). This may suggest that farther away locations recover more quickly after the shooting, or even that a substitution mechanism is at play. That is, customers now prefer farther away locations rather than those in close proximity to the site of the incident.
@@
We then compute an index measure of weekly aggregate visits for all non-truncated, non-spree shooting events as a preliminary exercise. The index takes a value of 100 for the week prior to the shooting (week -1) and is calculated either for all POIs up to 5 miles away from the shooting sites or separately by distance band. Distance bands are concentric zones surrounding a shooting site, with cutoffs determined by the quantiles of the distances between POIs and the shooting site. For instance, the first quartile distance band comprises POIs located within the bottom 25% of the distances, while the second quartile distance band includes POIs located between the 25th and 50th percentiles of the distances. Distance quantiles are computed at the shooting event level, and thus the sizes of the distance bands can differ from place to place. We prefer quantiles over raw distance (miles) to define our distance bands since it accounts for differences in POI density in the various shooting areas and thus captures “perceived distance” in a local community. [^9] A potential concern is that if POIs are highly concentrated in a specific area, the distance bands based on quantiles can be physically too close, failing to represent distinct areas in a community. Figures C.1 and C.2 in online Appendix C demonstrate that this is not a concern in our data. In these figures, we plot the relationship between raw distance and distance quantile, as well as the density of a given POI type at each distance quantile. Overall, there appears to be a linear relationship between raw distance and distance quantile, implying POIs in our data are in general evenly distributed over space. POI types are also quite uniformly distributed.
@@
Figure 1(b)(i) displays a substantial reduction in aggregate visits beginning the week of the shooting, where recovery is not complete even after 20 weeks. Figure 1(c)(i) shows a similar decline in aggregate visits across all distance bands immediately following a mass shooting, but a faster recovery is observed for the POIs in the third and fourth quartile distance bands. This aligns with the patterns observed in Figure 1(a). Visits to the POIs in the third and fourth quartile distance bands rise even above the pre-shooting levels, suggesting a potential substitution effect.
@@
Figures 1(b)(ii) and 1(c)(ii) plot a similar index for active POIs, which displays the evolution of the number of POIs for which nonzero visits are reported in the data. The time-evolution of this index is much more subtle, indicating that the changes in mobility patterns are neither driven by POI closures, nor merely by missing some of the POIs during the period of analysis.
@@ 3 Empirical Strategy
Our empirical strategy for identifying the causal impact of mass shootings on community mobility and relocation of economic activities hinges on the assumption that there is no fundamental change in community mobility prior to the week of mass shooting. We argue that this is a reasonable assumption given that our analysis focuses on non-targeted and unpredictable shootings, as discussed in Section 2.1. Also, the pre-trends observed in the graphs in this section do not show any notable changes near the week of mass shooting, supporting the assumption.
@@
$\quad y_{i,t}$ $=$ $\sum_{b\in\mathcal{B}}\sum_{\ell=-10}^{\ell=-2}\beta_{b,\ell}D_{i,t}^{b,\ell}+\sum_{b\in\mathcal{B}}\sum_{\ell=0}^{\ell=20}\beta_{b,\ell}D_{i,t}^{b,\ell}+\delta_{Y(ST_{i})\times b}+\gamma_{M(ST_{i})\times b}+\epsilon_{i,t},$ (1)
where $y_{i,t}$ is weekly visits to POI $i$ in week $t$, the set $\mathcal{B}$ includes distance bands, [^10] See Section 2.4 for the definition of our distance bands. In our analysis, we use either two-quantile or four-quantile distance bands. Online Appendix B.2 includes some nonparametric estimation results, demonstrating that our results are robust to the cutoffs used to define the distance bands. and $\ell$ is the week number relative to shooting omitting the week right before the shooting (baseline week). The dummy $D_{i,t}^{b,\ell}=\textbf{1}\{B_{i}=b,t-ST_{i}=\ell\}$, where $B_{i}$ is the distance band to which POI $i$ belongs, and $ST_{i}$ is the week of the mass shooting associated with POI $i$. Therefore, $D_{i,t}^{b,\ell}$ is an indicator for POI $i$ being in distance band $b$ and relative week $\ell$, and thus $\beta_{b,\ell}$ calculates the average change in visits in distance band $b$ in week $\ell$ compared to the band’s weekly visits in the week right before the shooting. The specification includes distance band-specific fixed effects for shooting year, $\delta_{Y(ST_{i})\times b}$, and month, $\gamma_{M(ST_{i})\times b}$, controlling for potential business cycles at higher-levels. The standard errors are clustered at the POI level. [^11] We considered POI fixed effects, instead of the band-specific shooting year and month fixed effects. The parameter estimates were very similar to those reported in this paper, but caused computation issues when clustering the standard errors at the POI level.
We also compute DiD-style estimates that compare the magnitude of the impact across different distance bands (e.g., Figure 2(b)) by reparameterizing the model (1) such that
$\quad y_{i,t}$ $=$ $\sum_{\ell\in\mathcal{G}}\beta_{\ell}D_{i,t}^{\ell}+\sum_{b\in\mathcal{B}/\{b^{\ast}\}}\sum_{\ell\in\mathcal{G}}\beta_{b,\ell}D_{i,t}^{b,\ell}+\delta_{Y(ST_{i})\times b}+\gamma_{M(ST_{i})\times b}+\epsilon_{i,t},$ (2)
where $\mathcal{G}=\{-10,-9,\ldots,20\}/\{-1\}$, $D_{i,t}^{\ell}=\textbf{1}\{t-ST_{i}=\ell\}$, $b^{\ast}$ is the reference band and thus $\beta_{b,\ell}$ in the model (2) calculates the impact relative to that on the reference.

### v2

@@ Abstract
Using foot traffic data for over 153,000 points of interest (POIs) near the sites of 42 mass shootings (2018-2022, U.S.), we evaluate the spatial-temporal impact of these tragic events on community mobility and relocation of economic activity. Visits to nearby POIs decrease, while farther away POIs experience increased foot traffic, implying that communities shift their activities away from the shooting sites. The impact is stronger when stronger trauma responses are expected. Our results suggest that mass shootings drive significant spatial-temporal displacements of economic activity and can lead to welfare losses due to distortions in optimal choices of time and location.
@@ 1 Introduction
Mass shootings have a profound effect on survivors, witnesses, and communities that experience them. The substantial loss of innocent lives is tragic on its own and clearly demonstrates that these events come at a great cost. However, the broader literature in economics, psychology, and beyond suggests that loss of life does not constitute the sole effect of mass shootings. Rather, it is accompanied by a variety of consequences on various fronts, including: worsening mental health for both direct victims and the surrounding community (Lowe and Galea, 2017; Sharkey and Shen, 2021; Wozniak et al., 2020); exacerbation of political polarization (Donohue et al., 2019; Luca et al., 2020; Yousaf, 2021); and decreases in school enrollment and performance (Beland and Kim, 2016; Levine and McKnight, 2024; Bharadwaj et al., 2021). [^1] For a recent survey of the literature on the causes and impacts of mass shootings, see Yousaf (2023).
@@
Although these tragic events are the subject of a growing number of papers, evaluations of their economic consequences are rare. Brodeur and Yousaf (2022) is one of the few published studies that provide a comprehensive analysis of mass shootings from an economic perspective. Using annual data, the authors find that mass shootings reduce the number of jobs and establishments in targeted counties by about 1.8% and 1.3%, respectively. They also find detrimental effects on earnings and housing prices in the years following the shootings. [^2] Another important recent study, Cabral et al. (2025), analyzes 32 non-mass shootings at public schools in Texas from 1995 to 2016, and finds exposure to a shooting harms human capital accumulation and leads to worse employment rates and earnings in the long term.
@@
In contrast to Brodeur and Yousaf (2022), which presents year-over-year and county-level effects, our analysis is centered on a more immediate and direct economic outcome: changes in visits to locations within five miles of a shooting site, right after the incident. Specifically, this paper answers the following questions: Do mass shootings result in the displacement of economic activity, especially toward locations farther from the shooting sites? Do mobility changes differ by location characteristics, such as substitutability? And are these changes related to the perceived risk of victimization or traumatic responses to the event? By answering these questions, our paper uncovers important mechanisms underlying the long-term, macroeconomic effects of mass shootings documented in Brodeur and Yousaf (2022).
@@
For our analysis, we use the mass shootings that occurred in the United States between 2018 and 2022, as documented by Mother Jones. This list excludes shootings from more conventionally motivated crimes, such as robbery or gang activities, and thus allows our analysis to be grounded on non-targeted and unpredictable events. To measure mobility, we use Safegraph’s “weekly patterns” dataset, which includes foot traffic data and relevant characteristics for points of interest (POIs). Then, we employ event study and Difference-in-Differences (DiD) regressions to estimate the causal impact of mass shootings on community mobility over a period spanning ten weeks before and twenty weeks after the incidents. [^3] Vachuska and Movahed (2025) employ a similar approach to estimate the effect of gun violence on community mobility. Their analysis, however, considers all gun violence incidents, including robberies and domestic violence, and uses foot traffic data aggregated for census tracts over a much shorter time period (28 days before and after a shooting).
@@
Our empirical findings can be summarized as follows. First, visits to nearby POIs decrease significantly after a mass shooting, while farther away POIs experience increased foot traffic a few weeks later, suggesting spatial and temporal displacements of economic activity. For instance, in week 5 after a mass shooting, nearby POIs exhibit an average decline in visits of 26% relative to the week before the shooting, while farther-away POIs show a 13% increase in week 12. Second, POIs that can be substituted with relative ease due to their business nature or the presence of a large number of competitors in the region exhibit more intensive changes. POIs that are difficult or undesirable to switch in the short run, such as healthcare facilities, show no sizable effect. [^4] We discuss the welfare implications of the reallocation of economic activity in Section 4.1. Even without measurable changes, welfare can still be affected. If the preference for the current location decreases due to a mass shooting but relocation is infeasible because of limited alternatives, a welfare loss may occur. Finally, the mobility changes and relocation of economic activity are stronger when the shooting occurs in a previously safer area, when it receives more media coverage, and for POIs that are more heavily visited before the incident. Combined, the results suggest that people reallocate their activities across time and space due to these events, especially under the conditions where stronger trauma responses are expected.
@@
This paper contributes to the literature in several ways. First, it sheds light on an important driver for the long-term, macroeconomic effects of mass shootings found in other studies (e.g., Brodeur and Yousaf, 2022). After an event of this type, communities experience significant spatial-temporal displacements of economic activity, likely incurring friction costs or efficiency losses due to distortions in optimal location and timing decisions. This can ultimately worsen various macro-level outcomes, including wages, employment rates, and housing prices. Second, our results reveal behavioral changes in response to trauma, linking the economic consequences to those documented in studies on trauma psychology (Wozniak et al., 2020; Sharkey and Shen, 2021), the economics of crime (Braakman, 2012; Dustmann and Fasani, 2014), criminology and sociology (Garofalo, 1981; Hale, 1996), and public health (Marquet et al., 2020; Guite et al., 2006). Finally, this study finds that, in terms of per-POI impact, recreational spaces, places for temporary accommodation, educational facilities, especially primary and secondary, food and drink stores (e.g., grocery stores), restaurants, and bars are among the most impacted locations. This opens up an opportunity for future studies to use the variations in activities in these categories to identify other important relationships between economic variables.
@@ 2.2 Community Mobility
The foot traffic data used in this study comes from Safegraph’s weekly patterns dataset, which is based on a collection of cellphone location data from a large list of applications that send pings - that is, requests for GPS data - to the phone in which they are installed. [^6] Papers using the Safegraph data include Chen and Rohla (2018), Chen et al. (2022), Chen and Pope (2020). Safegraph aggregates pings to produce a large panel-type dataset containing the number of weekly visits to distinct points of interest (POIs). [^7] See https://www.safegraph.com/guides/visit-attribution-white-paper for more details on Safegraph’s methodology In the dataset, a point of interest is any physical place of a non-residential nature, such as a restaurant, a museum, or a church, and it can belong to any of 180+ industry groups. In addition to tracking the number of visits to individual POIs within specified date ranges, the data include information that allows us to identify the locations and types of POIs: unique ID, latitude and longitude, location name, and North American Industry Classification System (NAICS) code.
Our main analysis uses a total of 4.2 million observations for about 153 thousand distinct POIs located within five miles of the mass shooting sites, covering a period of 30 weeks for each event - 10 weeks before the shooting and 20 weeks after. Nonparametric estimation results, reported in Online Appendix B.2, support the choice of five-mile radius, where the impact of mass shootings on mobility vanishes beyond five miles from the shooting site.
Six shooting events took place in more than one location. These events, known as sprees, require a special handling since our analysis relies heavily on the distance between a POI and its associated shooting site. In these cases, an overall centroid may be used to represent the crime scenes, but we have opted to include all crime scenes as separate shooting sites, including all POIs within a five mile radius of each of the crime scenes, and assigning the minimum distance-to-shooting-site in the event of overlaps. This strategy improves precision and is also more adequate in our analysis given that, in some cases, the distance between crime scenes is substantial (e.g., the Atlanta Massage Parlor Shootings’ victims were 30 miles apart). In Online Appendix A, events are individually identified as being sprees or having truncated time series. [^8] Our Safegraph data start in January 2018 and end in December 2022, so the mobility data are truncated for some shooting events. For instance, the Marjorie Stoneman Douglas High School shooting, which occurred on February 14th, 2018, does not have a full set of 10 weeks of data before the event, and it is thus truncated on the left. Seven out of the 42 mass shootings have such truncated time series.
@@ 2.4 Descriptive Analysis
(b) Evolution of Aggregate Visits and Active POIs - By Distance Band
@@
*Note: Subfigure (b) plots the evolution of visits and active POIs, using the week immediately before the shooting as the baseline. Sprees and events with incomplete time series are excluded.
@@
This section offers an introductory description of the mobility patterns observed in the data. Figure 1(a) shows the changes in weekly visits to McDonald’s locations in Chicago after the Mercy Hospital shooting on November 19th, 2018.101010This exercise uses a franchise to minimize heterogeneity across locations. The y-axis represents the changes in weekly visits at each location compared to the week right before the shooting. Clearly, most McDonald’s locations experienced a decrease in weekly visits in both week 1 and week 15 after the shooting (red dots), but we do see some locations farther from the shooting site experienced an increase in visits in the latter period (green dots). This may suggest that farther away locations recover more quickly after the shooting, or even that a substitution mechanism is at play. That is, customers now prefer farther away locations rather than those in close proximity to the site of the incident.
@@
We then compute an index measure of weekly aggregate visits for all non-truncated, non-spree shooting events as a preliminary exercise. The index takes a value of 100 for the week prior to the shooting (week -1) and is calculated separately by distance band. Distance bands are concentric zones surrounding a shooting site, with cutoffs determined by the quantiles of the distances between POIs and the shooting site. For instance, the first quartile distance band comprises POIs located within the bottom 25% of the distances, while the second quartile distance band includes POIs located between the 25th and 50th percentiles of the distances. Distance quantiles are computed at the shooting event level, and thus the physical sizes of the distance bands can differ from place to place. [^11] The median POI distance for each shooting event is reported in Online Appendix A. The full relationship between quantile distance and physical distance (in miles) for each shooting event is shown in Figure C.1 in Online Appendix C.
We prefer quantiles over raw physical distance (in miles) to define our distance bands since it accounts for differences in POI density in the various shooting areas and thus captures relative, locally “perceived distance.” This measure of perceived proximity can reflect relative accessibility in a local community, and therefore facilitates linking the observed mobility changes in our results to the behavioral changes or potential welfare losses among community members. [^12] A potential concern is that if POIs are highly concentrated in a specific area, the distance bands based on quantiles can be physically too close, failing to represent distinct areas in a community. Figures C.1 and C.2 in Online Appendix C demonstrate that this is not a concern in our data. In these figures, we plot the relationship between raw distance and distance quantile, as well as the density of a given POI category at each distance quantile. Overall, POIs in the dataset are well spread across the areas surrounding the shooting sites, and POI categories are also relatively uniformly distributed.
@@
Figure 1(b) (i) displays a substantial reduction in aggregate visits beginning the week of the shooting, across all distance bands. A faster recovery is observed for the POIs in the third and fourth quartile distance bands, and recovery is not complete for the first and second quartile distance bands even after 20 weeks. This aligns with the patterns observed in Figure 1(a). Visits to the POIs in the third and fourth quartile distance bands rise even above the pre-shooting levels, suggesting potential spatial reallocation. Another interesting pattern is that the POIs in the second distance band exhibit a larger decline in weekly visits than those in the first distance band. To some extent, this may be due to the presence of authorities, visitors, or the media near the shooting sites.
@@
Figure 1(b) (ii) plots a similar index for active POIs, which displays the evolution of the number of POIs for which nonzero visits are reported in the data. The time-evolution of this index is much more subtle, indicating that the changes in mobility patterns are neither driven by POI closures, nor merely by missing some of the POIs during the period of analysis.
@@ 3 Empirical Strategy
Our empirical strategy for identifying the causal impact of mass shootings on community mobility and relocation of economic activity hinges on the assumption that there is no fundamental change in community mobility prior to the week of mass shooting. We argue that this is a reasonable assumption given that our analysis focuses on non-targeted and unpredictable shootings, as discussed in Section 2.1. Also, the pre-trends observed in Figure 1(b) do not show any notable changes near the week of mass shooting, supporting the assumption.
@@
$\quad y_{i,t}$ $=$ $\sum_{b\in\mathcal{B}}\sum_{\ell=-10}^{\ell=-2}\beta_{b,\ell}D_{i,t}^{b,\ell}+\sum_{b\in\mathcal{B}}\sum_{\ell=0}^{\ell=19}\beta_{b,\ell}D_{i,t}^{b,\ell}+\delta_{i}+\gamma_{t}+\epsilon_{i,t},$ (1)
where $y_{i,t}$ is weekly visits to POI $i$ in year-week $t$, the set $\mathcal{B}$ includes distance bands, [^13] See Section 2.4 for the definition of our distance bands. In our analysis, we use either two-quantile or four-quantile distance bands. Online Appendix B.2 reports results estimating the impact on mobility as a function of continuous distance. The overall patterns are consistent with those based on distance bands. and $\ell$ is the week number relative to shooting omitting the week right before the shooting (baseline week). The dummy $D_{i,t}^{b,\ell}=\textbf{1}\{B_{i}=b,t-ST_{i}=\ell\}$, where $B_{i}$ is the distance band to which POI $i$ belongs, and $ST_{i}$ is the year-week of the mass shooting associated with POI $i$. Therefore, $D_{i,t}^{b,\ell}$ is an indicator for POI $i$ being in distance band $b$ and relative week $\ell$, and thus $\beta_{b,\ell}$ calculates the average change in visits in distance band $b$ in week $\ell$ relative to the week right before the shooting. The specification includes POI fixed effects, $\delta_{i}$, and year-by-week fixed effects, $\gamma_{t}$, controlling for potential business cycles at local-levels. [^14] Due to the large number of POIs in the data (approximately 153,000 distinct locations), accounting for POI fixed effects while estimating Equation (1) is computationally challenging. In our case, it created some computation issues. Therefore, to facilitate estimation, we demean weekly visits by POI and estimate Model (1) without including POI fixed effects. We also considered alternative specifications with coarser time fixed effects, such as year–by-month effects, and the results were qualitatively unchanged. The standard errors are clustered at the POI level.
We also compute DiD-style estimates that compare the magnitude of the impact across different distance bands by reparameterizing Model (1) such that
$\quad y_{i,t}$ $=$ $\sum_{\ell\in\mathcal{G}}\beta_{\ell}D_{i,t}^{\ell}+\sum_{b\in\mathcal{B}/\{b^{\ast}\}}\sum_{\ell\in\mathcal{G}}\beta_{b,\ell}D_{i,t}^{b,\ell}+\delta_{i}+\gamma_{t}+\epsilon_{i,t},$ (2)
where $\mathcal{G}=\{-10,-9,\ldots,19\}/\{-1\}$, $D_{i,t}^{\ell}=\textbf{1}\{t-ST_{i}=\ell\}$, $b^{\ast}$ is the reference band and thus $\beta_{b,\ell}$ in Model (2) calculates the impact relative to that on the reference.